OpenStudy (anonymous):

solve for x: X^(squarert(logx))=10^(2root2) nicer equation in comment below!

OpenStudy (anonymous):

\[x^\sqrt{logx}=10^{2\sqrt{2}} \]

OpenStudy (anonymous):

@kirbykirby

OpenStudy (kirbykirby):

is your log in base 10 or base \(e\) ?

OpenStudy (anonymous):

base 10

OpenStudy (kirbykirby):

Well regardless... you can try taking the log of both sides: \[\log\left( x^{\sqrt{\log x}}\right)=\log \left( 10^{2\sqrt{2}}\right)\\ \sqrt{\log x}\cdot \log x=2\sqrt{2}\log 10, ~~\text{ using the rule} \log a^b=b \log a\\ \left( \log x\right)^{1/2}\cdot \log x=2\sqrt{2} \log 10 \\ \left(\log x\right)^{3/2}=2\sqrt{2}\log 10, ~~\text{ using the law of exponents } a^xa^y=a^{x+y}\\ \left[\left(\log x\right)^{3/2}\right]^{2/3}=\left[2\sqrt{2}\log 10 \right]^{2/3}\\ \log x = \left[2\sqrt{2}\log 10 \right]^{2/3}\] Now here's the part where the base of the log matters... if it's base 10: then by the definition of the log, you get: \[ x=10^{\left[2\sqrt{2}\log 10 \right]^{2/3}}=10^2=100\] If it's in base \(e\), you get : \[ x=e^{\left[2\sqrt{2}\log 10 \right]^{2/3}}\approx 32.7022\]

OpenStudy (kirbykirby):

Ah ok lol sorry I didn't see your post above^ ignore the base \(e\) part :)

OpenStudy (kirbykirby):

in fact since it's in base 10, you can simplify from the beginning that \(\log 10 =1\) to simplify things a little bit

OpenStudy (anonymous):

lol the answer is just 10^2. i think you are supposed to do this without a gdc?

OpenStudy (kirbykirby):

gdc?

OpenStudy (kirbykirby):

and \(10^2=100\) as I showed above

OpenStudy (kirbykirby):

ah you mean graphical display calculator? (Googled it lol)

OpenStudy (anonymous):

oh yes!! haha sorry i didnt see that although i magnified the math notation to like 200 (im blind D:)

OpenStudy (kirbykirby):

ah lol sorry ya it's a bit small with all the exponents.. hm but to simplify it in more detail:

OpenStudy (anonymous):

why couldnt you simplify the log10 to 1? since it is base 10

OpenStudy (kirbykirby):

u just use exponent rules..\[x=\large 10^{\left[2\sqrt{2}\log 10 \right]^{2/3}}=\large 10^{\left[2\sqrt{2}\right]^{2/3}}=\large10^{2^{(2/3)}2^{1/2*2/3}}=10^{2^{(2/3)+(1/3)}}10^2=100 \]

OpenStudy (kirbykirby):

I just wrote it in general because I wasn't sure if it was base 10 or base e (I started writing before you replied xD), but I mentioned in my comment afterward, that you could have simplified it from the beginning since it was in base 10

OpenStudy (anonymous):

yep i saw that! thank you so much!!!